# Arecibo Wow! I: An Astrophysical Explanation for the Wow! Signal

### v1

@@ 8 Summary and conclusions
We observed narrow-band signals ($\Delta\nu\leq$ 10 kHz) similar to the Wow! Signal near the hydrogen line, though significantly weaker. These signals align with those originating from cold hydrogen (HI) cloud regions positioned in front of a galactic hydrogen background. We propose that the Wow! Signal was caused by the abrupt brightening of the hydrogen line in these clouds due to stimulated emission from a transient and intense radiation source, such as a magnetar flare or a soft gamma repeater (SGR).
The emission source must be positioned behind the cloud to facilitate observation of stimulated emission. This spatial alignment constitutes a rare occurrence and elucidates the previous lack of detection. In contrast, this inference suggests that the progenitor of the Wow! Signal resides at a greater distance than the clouds that produce the signal. Given the detectability of the clouds, as demonstrated in our data, this insight could enable precise localization of the signal’s origin and permit continuous monitoring for subsequent events.
Our hypothesis accounts for all observed characteristics of the Wow! Signal, introduces another source of false positives in technosignature searches, and suggests that this signal represented the first recorded astronomical maser flare in the hydrogen line.
We will continue with Arecibo Wow! exploring our extensive sets of observations from Arecibo REDS. Future studies will incorporate archival data from the Arecibo Observatory, which is now hosted at the Texas Advanced Computing Center (TACC).
@@
We extend our gratitude to former staff members of the Arecibo Observatory, Tapasi Ghosh and Christopher Salter, for helping us with our observations and for their valuable comments and suggestions. We are grateful to Anish Roshi Damodaran (UCF) for his assistance with data calibration. We finally thank Edgard Rivera-Valentín (Johns Hopkins University Applied Physics Laboratory) and Robert F. Minchin (National Radio Astronomy Observatory) for telling us that the Arecibo Observatory was easy to use, not true, but too late for that.

### v2

@@ 8 Summary and conclusions
We observed from the Arecibo Observatory narrowband signals ($\Delta\nu\leq$ 10 kHz) similar to the Wow! Signal near the hydrogen line, though significantly weaker. These signals are attributed to small cold Hi cloud regions within the galactic ISM. We propose that the Wow! Signal was caused by the abrupt brightening of the hydrogen line of these clouds triggered by an intense radiation source, such as a magnetar flare or a soft gamma repeater (SGR).
The brightening of the cloud is similar to that from a maser flare but short-lived, from seconds to minutes. Astrophysical superradiance and masers involve coherent radiation, but they differ fundamentally in their mechanisms. Superradiance tends to produce short, intense bursts, whereas masers are associated with continuous or pulsed emission and are more commonly observed in various astrophysical environments. In particular, we showed that Dicke’s supperradiance can explain the intensity and duration of the signal. The Wow! Signal might signify the first documented instance of an astrophysical maser-like flare or supperradiance of the 21-cm atomic hydrogen line.
Our hypothesis explains all the discernible attributes of the Wow! Signal, namely its narrowband frequency, duration, intensity, absence of modulation, and rarity. However, it does not conclusively rule out the possibility that the signal may have originated from alternative sources, encompassing both artificial terrestrial or extraterrestrial origins.
We advocate for a thorough examination of archival data as well as new observational efforts to accurately pinpoint the provenance of the signal or identify analogous wow signals events. Since these Hi clouds are easy to recognize, it may be possible to determine the exact location of the source of the Wow! Signal. Identifying the trigger source, however, would prove challenging. It may be close or behind the cloud location, or far in the background.
These signals are likely to escape detection, not only due to their scarcity and association with small Hi clouds, but also because contemporary observational strategies typically focus on broadband signals, prolonged integration times, or target much narrower signal bandwidths. It was not until we engaged in the search for 10 kHz narrowband signals, utilizing a drift (or mapping) mode, that the small Hi clouds came to our attention as a possible source for the Wow! Signal.
Our study did not conclude that the Wow! Signal constituted evidence of a signal emanating from an extraterrestrial civilization. However, null results are instrumental in refining future technosignature searches. NASA and other agencies remain committed to their search for biosignatures within the atmospheres of potentially habitable exoplanets (Schwieterman et al., 2018).
The next question after a positive biosignature detection would be whether it is also associated with intelligent life. It would be unsatisfactory to claim ignorance due to the absence of interest on technosignature searches. Therefore, we also emphasize the need for more sustained, long-term technosignature searches, particularly near the hydrogen line. Prospective extraterrestrial civilizations may similarly acknowledge the scarcity of natural Hi masers and could potentially exploit them to garner attention. However, it is plausible that they also deem it rude to direct a maser toward another entity.
Our findings highlight once again the importance of reanalyzing historical datasets with different or more advanced methods, which could potentially uncover subtleties missed in previous analyses. For example, fast radio bursts (FRBs) were discovered in archival data and later confirmed by observations (Lorimer et al., 2007). In our study, the replication of the observation and analysis methods used by Ohio SETI was the key to the association with Hi clouds.
Our Arecibo Wow! project will continue examining the rest of our datasets, which consist of numerous hours of observing stars with planets at frequencies ranging from 1 to 10 GHz. We will expand our narrowband search to incorporate artificially dispersed and modulated wideband signals, among other techniques (Li et al., 2020). Subsequently, we will try to move the search to over fifty decades of data (3 petabytes) from the Arecibo Observatory, now archived and available to the scientific community at the Texas Advanced Computing Center (TACC). We also anticipate the dissemination of calibrated and reduced datasets of Arecibo Wow! to the broader scientific community.
The Arecibo Observatory is currently a center for education and research in science, technology, engineering, and mathematics (STEM) known as the Arecibo C3. Looking to the far future, the facility may potentially resume its role in astronomical observations through initiatives such as the Next Generation Arecibo Telescope (nGAT) or participation as a site for the Next Generation Very Large Array (ngVLA) (Anish Roshi et al., 2021; Wilner et al., 2024). In the near term, a reestablishment of its 2.5 to 14 GHz 12-meter telescope should provide new opportunities for astronomical research and education (Roshi et al., 2024).
Acknowledgments
@@
We extend our gratitude to former staff members of the Arecibo Observatory, Tapasi Ghosh and Christopher Salter, for helping us with our observation protocols and for their valuable comments and suggestions. We are grateful to Anish Roshi Damodaran (UCF) for his assistance with data calibration and comments. We thank Bob Dixon (formerly Big Ear staff), Robert F. Minchin (NRAO), Jean-Luc Margot (UCLA), Joshua Peek (APL), and Joel Weisberg (Carleton College) for valuable comments on the manuscript.
We also thank the computational resources provided by the High Performance Computing Facility (HPCf) of the University of Puerto Rico (UPR), Juan Padilla (SoftwareOne), Amazon Web Services (AWS), and all the people who contributed to save our Arecibo Observatory data. Finally, we thank Edgard Rivera-Valentín (JHU-APL) for encouraging us to use the Arecibo Observatory because it was easy to use; not true, but too late for that.
Arecibo Observatory, Big Ear